# active distribution network DER distributed energy resource

### v2

@@ III-B Benchmark Distribution System
The IEEE33 benchmark feeder containing $1.4MWp$ of PV injections and $2.7MWp$ load is selected as a benchmark system. Four controllable DERs are considered, in this case all BESSs with a total energy capacity of $24MWh$ and a total power capacity of $6MW$. The BESS capacities were obtained by increasing the storage capacity until the self-dispatching version of the problem became feasible, allowing a comparison between the different aggregation versions. Time-coupled uncertainty sets are used for both PV and load. Specifically, the uncertainty sets obtained by considering the empirical $90\%$ coverage were used. Figure 4(a) shows the results when only the base power trajectory is controlled in a stacked plot. In this case more flexibility can be offered at the transmission level as uncertainties only impact potential congestions and other grid constraints but do not reserve energy or power capacity of the flexible DERs at the distribution level. aFRR is selected due to the higher service prices in this case, with no FCR provision for both the baseload and self-dispatching due to the lower prices.
@@
Figure 5 shows the results for the self-dispatching case, in which the flexibility is first used to compensate local uncertainties before services are offered to the upper-level system. Only upward aFRR is offered due to the limited flexibility. Note that the GCP baseload power shown in Figure 5(a) does not represent the total local balancing energy budget as both upwards and downwards capacity are reserved. Much less flexibility is available in the self-dispatching case, compared to the baseload case, as a large energy budget is reserved to balance deviations of the stochastic prosumption from the baseload. To show the utilization of the DERs storage assets and validate the aggregation sets, we run MC simulations, where for each selected set of flexibility actuations, we sample the prosumption from the considered uncertainty sets. The BESS SOEs trajectories are shown in Figure 6 for the self-dispatching case, showing the storage limits are satisfied. Additionally, note that the SOE stays far from the limits. This is due to the energy capacity reserved to balance the uncertain prosumption, for which only a limited number of realizations where used in the a posteriori sampling. Other quantities such as the ADN’s nodal voltages and branch currents are also validated as shown later in Section III-D.
@@ III-C Flexibility Potential in Realistic Distribution Systems
To demonstrate the multifeeder approach presented in Section II-F, we solve the aggregation problem for five similar ADNs obtained from the synthetic networks presented in [45], with the network IDs given in Table I. The loading and PV integration for all networks are added in Table I together with the controllable DERs characteristics. The considered voltage range at the slack node of the individual ADNs or feeders is set to $\delta=0.02\>pu$. Figure 7 shows the results for two cases of the transformer rating at the full network GCP. The service price of FCR was considered double the value reported in Figure 2(b) to show the provision of all services. Reducing the transformer rating to $27MW$ limits the available flexibility. For the symmetric FCR service, this leads to a reduction in both directions although the transformer constraint is only binding for additional power consumption.
@@
TABLE I: Characteristics of ADNs for the multifeeder case.
@@
For the application of the proposed aggregation approach, the computational complexity of the method is also relevant. Due to the complex nature of the problem, the computation time increases unacceptably when directly considering large distribution networks. The multifeeder approach allows to decompose large distribution networks per feeder to keep the problem tractable. For the two cases on the IEEE33 benchmark case, the computations times were 30 and 18 minutes for the self-dispatching and the baseload case respectively. For the multifeeder aggregation, the mean and max time per feeder were 24 and 45 minutes respectively. We recall that thanks to the robust voltage constraints, each feeder problem can be solved in parallel. The problem combining the feeders takes a few seconds. Solving the multifeeder problem in one go (without the disaggregation in feeders) was intractable on the adopted computer (a MacBook M1 Max with 32GB RAM).
III-D Importance of grid-aware flexibility aggregation.
@@
This section shows how the network constraints impact the aggregated flexibility. To this end, the aggregated flexibility for the IEEE33 network is compared with the case where all network constraints are neglected. To create additional congestions highlighting the importance of modeling the network constraints, we set all line ampacity limits to a third of their nominal value. For the baseload case, removing network constraints means uncertainty plays no role when considering only BESSs as flexible DERs as non of the constraints are affected by uncertainty drivers. Therefore the full flexibility can then be used for TSO requirements. Figure 8(a) shows the aggregated flexibility obtained with and without grid constraints. When the grid constraints are not accounted for, more flexibility can be offered at time steps with higher forecast service prices (cfr Figure 2(b)). Due to the time-coupling, more flexibility is offered in the grid-aware approach for some time steps. Note that the additional flexibility offered in the grid-unaware case is the result of an overestimation of the available flexibility and is not deliverable due to the violation of grid constraints. Also here, no FCR is provided. Figure 8(b) shows histograms of the line loading for the lines connecting the BESSs for the grid-aware and grid-unaware cases. Currents are computed a posteriori for a set of MC simulations where both the flexibility exchanged and the uncertain prosumption are varied. The frequent violations of the line constraints in the grid-unaware case show the importance of accounting for grid constraints when determining the available flexibility. The results also show that the exact currents slightly differ from the ones predicted by the linear model in the optimization. Therefore in some cases, the true currents may exceed the ampacity limits, as shown for example in Figure 8(b) for BESS2, but in general the limits hold.

### v3

@@ III-B Benchmark Distribution System
The IEEE33 benchmark feeder containing $1.3MWp$ of PV injections and $2.6MWp$ load is selected as a benchmark system. Four controllable DERs are considered, in this case all BESSs with a total energy capacity of $24MWh$ and a total power capacity of $8MW$. The BESS capacities are obtained by increasing the storage capacity until the self-dispatching version of the problem became feasible, allowing a comparison between the different aggregation versions. Time-coupled uncertainty sets are used for both PV and load. Specifically, the hyperbox uncertainty sets obtained by considering the empirical $90\%$ coverage are used. Figure 5(a) shows the results when only the base power trajectory is controlled in a stacked plot. In this case more flexibility can be offered at the transmission level as uncertainties only impact potential congestions and other grid constraints but do not reserve energy or power capacity of the flexible DERs at the distribution level. Only aFRR is selected due to the higher service prices in this case, with no FCR provision for both the baseload and self-dispatching due to the lower prices.
@@
Figure 6 shows the results for the self-dispatching case, in which the flexibility is first used to compensate local uncertainties before services are offered to the upper-level system. Only upward aFRR is offered due to the limited flexibility. Note that the GCP baseload power shown in Figure 6(a) does not represent the total local balancing energy budget as both upwards and downwards capacity are reserved. Much less flexibility is available in the self-dispatching case, compared to the baseload case, as a large energy budget is reserved to balance deviations of the stochastic prosumption from the baseload. To show the utilization of the DERs storage assets and validate the aggregation sets, we run MC simulations, where for the point in the aggregation set that maximizes the provided flexibility, we sample the load and weather conditions from the out-of-sample scenario set. Out of the 400 scenarios, 25 are outside of the considered uncertainty set ($6.25\%$), leading to 25 violations for the self-dispatching case. For the baseload case, these scenarios outside of the uncertainty set do not lead to violations. This occurs because the actuation of controllable DERs is more dependent on the uncertainty for the self-dispatching case and the resource constraints are the most stringent in this use case. Other quantities such as the ADN’s nodal voltages and branch currents are validated later in Section III-E.
@@ III-C Flexibility Potential in Realistic Distribution Systems
To demonstrate the multifeeder approach presented in Section II-F, we solve the aggregation problem for five similar ADNs obtained from the synthetic networks presented in [19], with the network IDs given in Table IV. The loading and PV integration for all networks are added in Table IV together with the controllable DERs characteristics. The considered voltage range at the slack node of the individual ADNs or feeders is set to $\delta=0.02\>pu$. Figure 7 shows the results for two cases of the transformer rating at the full network GCP. The service price of FCR was considered double the value reported in Figure 2(b) to show the provision of all services. Reducing the transformer rating to $27MW$ limits the available flexibility. For the symmetric FCR service, this leads to a reduction in both directions although the transformer constraint is only binding for additional power consumption.
@@
TABLE IV: Characteristics of ADNs for the multifeeder case.
For the application of the proposed aggregation approach, the computational complexity of the method is also relevant. Due to the complex nature of the problem, the computation time increases unacceptably when directly considering large distribution networks. The computational complexity of the problem is mostly due to the large number of conic constraints, required to reformulate the robust problem. The number of conic constraints in this problem can be expressed in (29) as a function of the key problem dimensions. The first one is the number of timesteps $T$. The other problem parameters affecting the complexity are the numer of considered services $n_{s}$, the number of uncertainty sets $n_{u}$, the number of nodes $N_{n}$ and lines $N_{l}$, the number of controllable DERs and storage assets $N_{s}$.
$n_{cones}=(n_{s}+n_{u})*(N_{n}+N_{l}+2*N_{c}+N_{s})*2*T$ (29)
@@
The multifeeder approach allows to decompose large distribution networks per feeder to keep the problem tractable. For the two cases on the IEEE33 benchmark case, the computations times were 30 and 18 minutes for the self-dispatching and the baseload case respectively. For the multifeeder aggregation, the mean and max time per feeder were 24 and 45 minutes respectively. We recall that thanks to the robust voltage constraints, each feeder problem can be solved in parallel. The problem combining the feeders takes a few seconds. Solving the multifeeder problem in one go (without the disaggregation in feeders) was intractable on the adopted computer (a MacBook M1 Max with 32GB RAM).
III-D Extension to multiple DERs types
In this section, we show the impact of allowing PV installations and HPs to participate in the flexibility provision. The two first PV installations in the IEEE33 case, with power ratings increased to $500MWp$ and $800MWp$, are considered controllable, together with a HP with a maximum electric power consumption of $250kW$. Figure 8 shows which services are offered and how the different types of DERs participate in the flexibility provision. Consistent with the operation of the different DERs, the BESSs are allowed to provide all services, while the HPs do not participate in the FCR provision and the PV installations only participate in the aFRR provision and do not participate to the compensation of prosumption uncertainty.
(a) Aggregated flexibility offered at the feeder GCP.
(b) Contribution of BESSs to the offered flexibility.
(c) Contribution of other resources to the offered flexibility.
Figure 8: Aggregated flexibility for the combined DER case.
Figure 9 shows the evolution of the SOE of the BESSs and the temperature evolution for the HP buffer tanks for the 400 out-of-sample scenarios used to validate the selected flexibility aggregation set. This clearly shows that both the HP and PV installations participate in the flexibility provision. The PV installations only provide aFRR down, as curtailment is disallowed to avoid participation in the compensating of the prosumption uncertainty. The HPs participate both to the baseload control and the aFRR provision. This also allows the BESSs to increase their aFRR provision. The out-of-sample analysis led to constraint violations in $2.75\%$ of the cases, with $11.25\%$ of the cases outside of the uncertainty set.
(a) HP temperature for the out-of-sample scenarios.
(b) BESSs SOE for the out-of-sample scenarios.
Figure 9: Evolution of the storage for the combined DER case.
III-E Importance of grid-aware flexibility aggregation.
@@ III-D Extension to multiple DERs types
This section shows how the network constraints impact the aggregated flexibility. To this end, the aggregated flexibility for the IEEE33 network is compared with the case where all network constraints are neglected. To create additional congestions highlighting the importance of modeling the network constraints, we set all line ampacity limits to a third of their nominal value. For the baseload case, removing network constraints means uncertainty plays no role when considering only BESSs as flexible DERs as none of the constraints are affected by uncertainty drivers. Therefore the full flexibility can then be used for TSO requirements. Figure 10(a) shows the aggregated flexibility obtained with and without grid constraints. When the grid constraints are not accounted for, more flexibility can be offered at time steps with higher forecast service prices (cfr Figure 2(b)). Due to the time-coupling, more flexibility is offered in the grid-aware approach for some time steps. Note that the additional flexibility offered in the grid-unaware case is the result of an overestimation of the available flexibility and is not deliverable due to the violation of grid constraints. Also here, no FCR is provided. Figure 10(b) shows histograms of the line loading for the lines connecting the BESSs for the grid-aware and grid-unaware cases. Currents are computed a posteriori for a set of MC simulations where both the flexibility exchanged and the uncertain prosumption are varied. The frequent violations of the line constraints in the grid-unaware case show the importance of accounting for grid constraints when determining the available flexibility. The results also show that the exact currents slightly differ from the ones predicted by the linear model in the optimization. Therefore in some cases, the true currents may exceed the ampacity limits, as shown for example in Figure 10(b) for BESS2, but in general the limits hold.